OpenStudy (anonymous):

Does the equation have a minimum or maximum value and what is it? Equation= y = 4x^2-6x-4 It's vertex: (3,-40)

OpenStudy (anonymous):

this equation has a minimum....because y'' is positive 1)take the derivative ....equate it to zero 2)solve for X 3)substitute the value of X in the original equation to get that value....

OpenStudy (anonymous):

Well, the equation has a min because the sign of the leading coefficient is positive. To find it, use x = -b/2a. Then use plug that x-value back into the equation to get y. So your minimum will be at (x,y). Ok?

OpenStudy (anonymous):

theprophet: whats the derivative? coolmath : how do i get the x alone ? i forgot how to do that.

OpenStudy (anonymous):

Y = ax^2 + bx + c is the form your equation is in. b = -6, a = 4.

OpenStudy (anonymous):

So, x = -(-6)/8. Or x = 3/4

OpenStudy (anonymous):

Plug that back into your equation to get your y-value.

OpenStudy (anonymous):

fine fine!!! u can try perfect square method where u make the equation into a perfect square plus sum constant....